OpenStudy (anonymous):

72+4-14d=36 i need to show how i got the answer

OpenStudy (anonymous):

So, you have 4 terms and you are trying to solve for "d". How do you get terms from one side of an equation to the other? Hint: when x=y, then a + x = a + y. Adding the same number to both sides preserves the equality.

OpenStudy (anonymous):

ok so if 72+4=76-14d=36 then d=40 right

OpenStudy (anonymous):

And before even trying to move terms from one side to the other, a more fundamental consideration is that you want to isolate the term with the "d" in it. That's why you even want to consider moving terms from one side to the other.

OpenStudy (anonymous):

Get d on the left-hand side by itself because there are a couple more steps. "d' is not 40. But you are on the right track with 76-14d=36 . Get 76 over to the right first.

OpenStudy (anonymous):

d=76

OpenStudy (anonymous):

can u just tell me the awnser lol

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

Hint: 76-14d=36 -> -76 + 76-14d=36 - 76 -> (-76 + 76) -14d=36 - 76 Study this sequence a little bit and then simplify the last equation. Go through all the above steps FIRST. You will get it, but you have to work it. I could tell you the answer, but then I would be doing you absolutely no good by doing that. The idea is for you to figure out how to do these all on your own eventually. Don't try to take shortcuts in your learning, especially of math. Math is extremely vertical. If you don't take the time to learn pre-algebra, you'll never learn algebra or trig or calculus. Your learning and growth will suffer, so work hard, it's definitely worth it, and in ways I can't explain to you right now.

OpenStudy (anonymous):

thanks i got it u helped alot

OpenStudy (anonymous):

Parentheses help a lot when all else looks bleak. I wish you well.